OpenStudy (anonymous):

Help plz will give medal who helps! 5(-3x - 2) - (x - 3) = -4(4x + 5) + 13

OpenStudy (texaschic101):

5(-3x - 2) - (x - 3) = -4(4x + 5) + 13 --- distribute through the parenthesis -15x - 10 -x + 3 = -16x - 20 + 13 -- combine like terms -16x - 7 = -16x - 7 -16x + 16x = -7 + 7 0 = 0 INFINITE SOLUTIONS